HanAkoSolo (jamierox4ev3r):

Solve:

HanAkoSolo (jamierox4ev3r):

Give me a second to type the equation, thanks :)

HanAkoSolo (jamierox4ev3r):

Solve: \[2\tan ^{2}\theta \sin \theta - \tan ^{2}\theta=0\] for all values in [0, 360] <---degrees

HanAkoSolo (jamierox4ev3r):

@ganeshie8 any ideas?

ganeshie8 (ganeshie8):

it seems \(\tan^2\theta\) is common in both the terms you may start by factoring it out

HanAkoSolo (jamierox4ev3r):

Alright. So that would be \( \tan( 2\tan \theta \sin \theta-\tan \theta )\) right?

ganeshie8 (ganeshie8):

pull out the entire \(\tan^2\theta\)

HanAkoSolo (jamierox4ev3r):

the entire thing? then that would leave you with.. just sine I believe. So \(\tan^{2} \theta (2\sin \theta-\theta)\)

ganeshie8 (ganeshie8):

wait a sec, \(\tan^{2} \theta (2\sin \theta-\color{Red}{\theta})\) are you really sure that \(\color{Red}{\theta}\) stays back there ?

HanAkoSolo (jamierox4ev3r):

wait... no it wouldn't. whoops xD

ganeshie8 (ganeshie8):

\(\tan^2\theta\) is ONE SINGLE thing \(\tan^2\) and \(\theta\) are not two pieces

HanAkoSolo (jamierox4ev3r):

^ yes I recall that now. But wouldn't something have to stay there?

HanAkoSolo (jamierox4ev3r):

oh wait. wouldn't it just be -1?

ganeshie8 (ganeshie8):

good, please fix it also what happened to \(= 0\)

HanAkoSolo (jamierox4ev3r):

oh it's still there, I just lopped it off. So as of now, this is where we are. \(\tan^{2} \theta (2\sin \theta-1)=0\)

ganeshie8 (ganeshie8):

looks good! what are you going to do next

HanAkoSolo (jamierox4ev3r):

I'm not sure....seems to me like there is a trigonometric identity somewhere in there perhaps.

ganeshie8 (ganeshie8):

Easy, actually we're almost done! all we need to know is the "zero product property"

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

so far we have : \(\color{red}{\tan^{2} \theta} \color{blue}{(2\sin \theta-1)}=0\) By zero product property, \(\color{red}{\tan^2\theta}=0\) or \(\color{blue}{2\sin\theta-1}=0\) solve each of them separately using unit circle

HanAkoSolo (jamierox4ev3r):

so we could set either one to zero. And solve for theta \((\theta)\)

ganeshie8 (ganeshie8):

Exactly! thats what zero product property tells us

HanAkoSolo (jamierox4ev3r):

wow. that actually makes so much sense. Alright, so let me solve each separately and I'll be back with you on what I got

ganeshie8 (ganeshie8):

take ur time you may have to use unit circle as you're asked to find all the solutions between 0 and 360

HanAkoSolo (jamierox4ev3r):

Noted.

HanAkoSolo (jamierox4ev3r):

Alright. So I got that \(\Large\sin \theta = \frac{1}{2}\) So from that, I found that \(\Large\theta = 30 (degrees), 150 (degrees)\)

HanAkoSolo (jamierox4ev3r):

I wasn't as sure for \(\tan \theta\), but I think that \(\Large\tan \theta =0\) and from that, wouldn't \(\Large\theta = 0 (degrees)\) ?? **not sure**

ganeshie8 (ganeshie8):

Excellent! \(\tan\theta=0\) should give you 3 solutions

ganeshie8 (ganeshie8):

il give a hint : \(\tan\theta = \tan(\theta+180)\)

HanAkoSolo (jamierox4ev3r):

oh wait i think i see. 90 degrees would be undefined

ganeshie8 (ganeshie8):

In other words : adding 180 to the angle wont change the value of \(\tan(\theta)\)

HanAkoSolo (jamierox4ev3r):

right.

HanAkoSolo (jamierox4ev3r):

I get that.

ganeshie8 (ganeshie8):

\(\tan\theta=0\) \(\implies \theta = \{0,\pm 180,\pm 360,\pm540,\ldots\}\) since you want the solutions between 0 and 360 you just pick \(0,180,360\)

HanAkoSolo (jamierox4ev3r):

oh and btw, I can include 0 degrees, but not 360. I was supposed to write the problem like [0,360) instead of [0,360]. So technically, for this problem, there are only two solutions that come from \(\tan \theta =0\)

ganeshie8 (ganeshie8):

Oh then you're absolutely right!

HanAkoSolo (jamierox4ev3r):

alright, so in total, all the solutions are 0, 30, 150, and 180! (all in degrees) oh my goodness, thank you so much! This review material was so distant in my memory, until you came along of course. Your patience is outstanding, thank you

ganeshie8 (ganeshie8):

np your patience is outstanding too! keep it up :)

HanAkoSolo (jamierox4ev3r):

d'aw you flatter me :')